Number Series
1. 75% is to 1/2 as 45% is to __________
2. 2 is to 50 as 3.2 is to _____________
3. 14 is to 28 as 5 is to __________
4. 1 is to 1/4 as 5/25 is to _________
5. 0.35 is to 7 as 0.45 is to ________

Answer by rfer(16322)

1) 30
2)
2/50=3.2/x
2x=160
x=80
3)
14/28=5/x
14x=140
x=10
4)
1/(1/4)=(5/25)/x
x=5/100
x=1/20
You try number 5.
Bob